### v1

@@ 5.1 Effective densities
The first step in calculating the kinetic energy and momentum density distributions is determining the effective density of the material sampled by each gas tracer across distance channels. In principle, CO is associated with densities roughly above 100 cm-3. Yet, the Edenhofer et al. (2024) 3D dust models do not have the spatial resolution to detail the regions where these densities are prevalent. Moreover, observations indicate that the emission from these molecules also extends into low-column-density regions (see, for example, Roman-Duval et al., 2016). Similar arguments also complicate the identification of the material representing Hi across distance channels in the 3D dust models. Thus, we used the morphological correlation provided by the HOG method to estimate an effective density related to the line emission from each gas tracer.
We employed the map of relative orientation angles between gradients for each distance and LOS-velocity channel pair. We split this map into 9 $\times$ 9 blocks, computed $V_{\rm d}$ for each block, and calculated the mass contained in the blocks with $V_{\rm d}$$>$ 1.0, which roughly corresponds to a 1-$\sigma$ significance for $V_{\rm d}$. Given the limited number of independent gradient vectors in each block, the threshold on $V_{\rm d}$ is necessarily less rigid than that we used for the whole tile. More stringent values lead to null values of the effective density for most of the distance channels. Currently, the HOG analysis is limited by the angular resolution in the gas observations and the 3D dust reconstruction, which ultimately define the number of independent gradient vectors used in the morphological comparison.
The effective density for either gas tracer is determined by dividing the total mass in the blocks with $V_{\rm d}$$>$ 0 by the volume of the distance channel. Explicitly, for each distance channel $q$ in the portion of the Galactic plane identified with the subindex $k$, we identified a velocity channel $p^{*}$ that has the highest $V_{\rm d}$ for the gas tracer $X$. For that distance channel, we compute the effective nucleon density,
@@
We note that the effective densities calculated using Eq. (3) are not the specific density of a gas tracer but a combination of its volume filling factor and a selection based on its correlation with the 3D dust density. The value of $n^{\rm eff}$ aims to acknowledge the fact that not all of the matter traced by the 3D dust is related to the motions traced by the gas.
The top-left panel of Fig. 14 shows the effective density distribution across distance cells. By construction, the effective densities for the gas tracers are below the total dust mean. While the mean density from all the dust is around 0.42 cm-3, the effective densities are around 0.19 and 0.09 cm-3 for Hi and CO, respectively. The lower value for CO reflects that tracer’s lower volume filling factor, which implies a smaller contribution than that from Hi to the kinetic energy and momentum for each volume element in our reconstruction.
@@ 5.2 Kinetic energy
We estimated the average kinetic energy from the streaming motions for each distance channel $q$ as
@@
where $X$ corresponds to the LOS velocity tracer, either Hi or CO, $(v^{R19}_{\rm LOS})_{kq}$ is the expected LOS velocity from the Reid et al. (2019) Galactic rotation model for the Galactic plane position $k$ and the distance channel $q$, and $(v^{X}_{\rm LOS})_{kp^{*}q}$ is the LOS velocity from the line emission channel $p^{*}$ with the highest morphological correlation ($V_{\rm d}$) with the distance channel $q$. We calculated the effective volumetric density,
@@
using Eq. (3), the proton mass, $m_{p}$, and $\mu$ is the mean nucleon mass factor, which for solar metallicity is 1.402 (Asplund et al., 2009; Draine, 2011).
The $E_{\rm K}$ distribution is shown in the top-right panel of Fig. 14. We found that the average $E_{\rm K}$ is around 0.68 and 0.18 eV/cm3 for Hi and CO, respectively. These estimates are comparable to the 0.22 eV/cm3 obtained for the turbulent kinetic energy in the local ISM, obtained with a mean nucleon density $n$$=$ 30 cm-3 and velocity $v$$=$ 1 km s-1, or mean values $n_{\rm H}\right>$$=$ 1 cm-3 and $\sigma_{v}$$=$ 5.5 km s-1 (Draine, 2011). Given that we are only sampling one component of the velocity field, the actual $E_{\rm K}$ value can be up to a factor of three higher.
The $E_{\rm K}$ standard deviations are around 2.22 and 0.55 eV/cm3 for Hi and CO, respectively. These values, roughly a factor of four above the mean, suggest significant fluctuations of this quantity within the 2.5-kpc-diameter ISM parcel considered in this paper. The maximum $E_{\rm K}$ values are around 100 and 30 eV/cm3 for Hi In general, the highest $E_{\rm K}$ are associated with the highest $n^{\rm eff}$, suggesting that density is the dominant factor in our reconstruction of the kinetic energy distribution.
Figure 15 shows the $E_{\rm K}$ maps obtained from the Hi and CO emission. We found no evident similarity in the $E_{\rm K}$ distribution in the two gas tracers. However, for both tracers the highest $E_{\rm K}$ are found in the regions at $d$$<$ 300 pc. The $E_{\rm K}$ radial profile, presented in the top panel of Fig. 16, shows values up to three times over the mean for regions outside $d$$\approx$ 300 pc and up to 20 times for regions inside that range.
There are two prominent peaks in the $E_{\rm K}$ radial profiles shown in Fig. 16. One at $d$$\approx$ 100 pc, which is roughly the radius of the Local Bubble (see, for example, Zucker et al., 2022). The second peak is centered around $d$$\approx$ 200 pc, and is more prominent in CO. The $E_{\rm K}$ peaks at $d$$<$ 300 pc do not seem to result from a concentration of high $V_{\rm d}$ due to a bias in the morphological correlation toward near distances, as discussed in App. C. However, the angular resolution in the 3D dust reconstruction may limit the mapping of $E_{\rm K}$ enhancements to further distances.
One potential explanation for the $E_{\rm K}$ peak at $d$$\approx$ 100 pc is the LSR location within the Local Bubble. We are observing the gas motions from a location near the center of that cavity so that the peak will be at a similar radius in all directions. This is less likely to be the case for material that is further away from us.
@@
The $E_{\rm K}$ azimuthal profile, obtained by adding the energy contributions along the line of sight, is presented in the bottom panel of Fig. 16, It shows peaks toward $l$$\approx$ 40 to 90 for Hi and CO. This range covers the general directions along the axis of two large-scale structures identified in 3D dust reconstructions: the Split (Lallement et al., 2019) and the Radcliffe wave (Alves et al., 2020), shown in Fig. 3. Additional overdensities are found toward the fourth quadrant, roughly between $l$$\approx$ 290 to 340. There is also a singular $E_{\rm K}$ outlier in CO toward $l$$\approx$ 175 that is related with the high $E_{\rm K}$ values for distance channels within $d$$\approx$ 200 pc and around $d$$\approx$ 400 pc toward that direction in Fig. 15.
@@ 5.3 Momentum density and momentum
The distribution of $(p^{X}_{\rm K})_{kq}$ is shown in the bottom-left panel of Fig. 14. The mean of the values is around 23.3 and 4.4 $\times$$10^{35}$ dyn s/pc3 for the Hi and CO emission, respectively. However, there are large excursions from these mean trends, characterized by standard deviations of roughly 5.2 and 1.28 $\times$$10^{37}$ dyn s. This implies that the local fluctuations are much more important than the mean values.
Still, the anisotropy in the density distribution and $v_{\rm LOS}$ means that not all the flows cancel out perfectly. Adding all the LOS momentum vectors through the studied volume, we obtain a net amount of momentum for Hi of 299.7 $\pm$ 0.35 $\times$$10^{43}$ dyn s and points toward $l$$=$ 299 .7 $\pm$ 85 .36. For CO, the net momentum is 57.0 $\pm$ 0.1 $\times$$10^{43}$ dyn s toward $l$$=$ 285 .3 $\pm$ 92 .0. Despite the significant standard deviation in the orientation angle, the general orientation is similar for the Hi and CO. We note that this quantity is biased toward the furthermost distance channels, which cover larger volumes, and it is most likely a feature of the reconstruction boundaries rather than an indication of a net mass flux through the volume within 1.25 kpc from the Sun.
Figure 17 shows the $p$ maps derived from the radial motions sampled by the two gas tracers. We chose the momentum density units to facilitate the comparison with the expected radial momentum input from a single supernova remnant (SNR), which has been identified as roughly between 3 $\times$$10^{42}$ and 3 $\times$$10^{43}$ dyn s in numerical simulations of one SNR expansion in an inhomogeneous density field for different ambient medium’s densities, metallicities (Martizzi et al., 2015). The peak momentum density in Hi, found toward $l$$\approx$ 25at $d$$\approx$ 100 pc, corresponds to a momentum around 0.43 $\times$$10^{43}$ dyn s. If that quantity of motion could be exclusively attributed to the input from SNRs, it would take just a few SNe transferring thermal to kinetic energy with an efficiency of a few percent to produce it.
@@ 5.4 Mass flow rates
[…]
@@
The azimuthal $\dot{M}$ profiles are presented in Fig 19. They correspond to the mean and the standard deviations of $\dot{M}$ along the LOS for Hi and CO. The mean mass flows, displayed in Fig 19, show net matter displacements of roughly couple $10^{-3}$ M/yr for lines of sight toward the first Galactic quadrant and 250 $<$$l$$<$ 360. This could be interpreted as residual velocity from the Galactic rotation if not because it has the opposite sign of the expected LOS motions indicated on the right-hand-side panel of Fig. 11. Further interpretation of these mass flows should consider that they may be a product of the LSR position and the boundaries of the 3D dust reconstruction.
@@ 6 Discussion
We found that roughly 72% and 45% of the 10$\times$ 10 distance channels in Fig. 5 show a significant morphological correlation, $\max(V_{\rm d})_{v}$$>$ 2.87, between the 3D dust and the Hi and the CO emission, respectively. The lower correlation with CO an be explained by the lower volume filling factor of that tracer (Kwan & Sanders, 1986; Kalberla & Kerp, 2009). Clouds traced by CO are expected to be smaller and more concentrated than Hi structures. Hence, CO has fewer independent CO gradients for the HOG morphological comparison on the same sky area and at the same common angular scale set by the derivative kernel size.
@@
The lack of correlation in the rest of the channels can be attributed to several factors. First, under-densities in local ISM volumes, such as found in the regions around 600 $<$$d$$<$ 800  pc toward $-45$$<$$l$$<$ 30 and 1000 $<$$d$$<$ 1200  pc toward $250$$<$$l$$<$ 290, to mention a couple of examples. Second, there is an absence of line emission of some LOS velocity ranges toward particular lines of sight. We previously mentioned the region around the superbubble GSH 238+00+09, where there is very little CO, as illustrated in the right panel of Fig. 6. Identifying the exact physical conditions that produce the local ISM under-densities and a lack of emission toward particular regions is beyond the scope of this work. However, the maps in Fig. 5 and Fig. 6 provide a global context for identifying the conditions toward specific regions and their implications for the ISM dynamics at smaller scales.
@@
There is a third global factor that reduces the correlation between the 3D dust and the line emission. Observations of other ISM tracers, such as gamma-rays and the [Cii] and OH line emission, reveal that not all of the ISM material is traced by Hi and CO emission (see, for example, Grenier et al., 2005; Langer et al., 2014; Pineda et al., 2013; Busch et al., 2021). Including this material in the distance-velocity correlation is not straightforward. In the case of [Cii], there is no map with enough angular resolution and coverage to obtain significant spatial correlations using HOG. In the case of OH, the emission distribution involves the complex physical conditions that lead to the emission from that radical, which calls for a separate dedicated study (see, for example, Rugel et al., 2018; Dawson et al., 2022). We acknowledge that those and other line tracers should be accounted for in future reconstructions of the local ISM, adding richness to the results we present in this paper.
Yet, our analysis has shown we have enough distance channels where the correlation between the dust distribution and the line emission is significant enough to identify prevalent LOS velocities for different portions of the local ISM, which we discuss in the following section.
@@ 6.2 Local ISM motions
We used the morphological correlation across distances and LOS velocities to assign a $v_{\rm LOS}$ to the distance channels in each of the 10$\times$ 10 regions in which we divided the studied volume. This assignment simplifies the velocity field by identifying a single value to ISM parcels with volumes between 102 to 105 pc3. Thus, our velocity reconstruction represents the local ISM dynamics on spatial scales of tens of parsecs in the plane of the sky and around a parsec in the LOS direction.
@@
The maps of reconstructed LOS velocities shown in Fig. 11 indicate that the local LOS velocity field carries the imprint of the Galactic rotation pattern in both tracers, as inferred from the comparison with the expected $v_{\rm LOS}$ in the Reid et al. (2019) Galactic rotation model, shown in Fig. 10. The large-scale pattern of Hi and CO emission is deeply influenced by Galactic rotation, as shown in figure B.2 of (HI4PI Collaboration et al., 2016) and figure 3 of (Dame et al., 2001). However, the peculiar velocities of MCs within 3 kpc cannot be explained purely by Galactic rotation (see, for example, Stark & Brand, 1989).
@@
Figure 12 shows that the distribution of LOS streaming motions roughly centered around zero, within the uncertainties imposed by the velocity channel widths in the input data. The standard deviations of the rotation-subtracted LOS velocities are 12.1 and 6.1 km s-1 for Hi and CO, respectively. Previous studies have identified similar deviations from Galactic rotation. Clemens (1985) used the Massachusetts-Stony Brook Galactic plane CO survey to identify streaming motions with an amplitude of around 5 km s-1 on length scales larger than $0.22R_{\rm 0}$, where $R_{\rm 0}$ is the radius of the solar orbit around the Galactic center, so that $0.22R_{\rm 0}$ corresponds to a scale of between 1.87 and 2.2 kpc for the two values considered in that work. Brand & Blitz (1993) reported streaming motions with mean values of 12 km s-1 in their modeling of the spectro-photometric distances for a sample Hii regions and the LOS velocities for their associated MCs. The authors found a quadrupolar LOS velocity pattern, qualitatively comparable to those in Fig. 11, and report that the molecular gas is streaming past the LSR at about 3.8 km s-1, which are within a factor two we obtained for CO in the Solar vicinity. Our reported velocity dispersions in Hi are also comparable to those found in The Hi Nearby Galaxy Survey (THINGS; Walter et al., 2008; Schmidt et al., 2016). However, the physical conditions in their sample are not in general equivalent to those in the solar vicinity.

### v2

@@ 5.1 Effective densities
In the standard, line-emission-based PPV approach, the density and LOS velocity of an object are obtained by isolating a line component, identifying its centroid, and integrating the emission (see, for example, Miville-Deschênes et al., 2017). Thus, the density and the LOS motion are defined for the same PPV object. Our analysis defines the object in PPD space as a slice of 3D dust density for a particular sky area. We assigned a $v_{\rm LOS}$ to that object using the HOG method through morphological matching. However, not all the 3D dust slice is represented by the $v_{\rm LOS}$ from one tracer. For example, the HI and CO emission can match different portions of a 3D dust slice, and each portion can have different velocities, as illustrated in Fig. 4. Thus, we introduce an “effective density” corresponding to the portions of the 3D dust slice with a significant morphological match with either gas tracer. This approach has the advantage of separating the Hi and the CO motions and assigning them a portion of the 3D dust instead of using the whole slice, which would lead to an overestimation of the energy and momentum densities.
We employed the map of relative orientation angles between gradients for each 3D density slice and $v_{\rm LOS}$-channel pair. We split this map into 9 $\times$ 9 blocks, computed $V_{\rm d}$ for each block, and calculated the mass contained in the blocks with $V_{\rm d}$$>$ 1.0, which roughly corresponds to a 1-$\sigma$ significance for $V_{\rm d}$. Given the limited number of independent gradient vectors in each block, the threshold on $V_{\rm d}$ is necessarily less rigid than that we used for the whole tile. Currently, the HOG analysis is limited by the angular resolution in the gas observations and the 3D dust reconstruction, which ultimately define the number of independent gradient vectors used in the morphological comparison. More stringent values lead to null values of the effective density for most of the distance channels.
The effective density in each density slice is determined by dividing the total mass in the blocks with $V_{\rm d}$$>$ 1.0 by the volume of the slice. Explicitly, for each distance channel $q$ in the portion of the Galactic plane identified with the subindex $k$, we identified a velocity channel $p^{*}$ that has the highest $V_{\rm d}$ for the gas tracer $X$. For that distance channel, we compute the effective nucleon density,
@@
The top-left panel of Fig. 14 shows the effective density distribution across distance cells. By construction, the effective densities for the gas tracers are below the total dust mean. While the mean density from all the dust is around 0.51 cm-3, the effective densities are around 0.27 and 0.14 cm-3 for Hi and CO, respectively. The lower value for CO reflects that tracer’s lower volume filling factor, which implies a smaller contribution than that from Hi to the kinetic energy and momentum for each volume element in our reconstruction.
We note that the effective densities calculated using Eq. (3) are not the specific density of a gas tracer but a combination of its volume filling factor and a selection based on its correlation with the 3D dust density. For example, CO is associated with densities roughly above 100 cm-3, but the Edenhofer et al. (2024) 3D dust models do not have the spatial resolution to detail the regions where these densities are prevalent. The value of $n^{\rm eff}$ does not aim to estimate the specific volumetric density of the gas but rather identify the portion of the 3D dust slice associated with the motion of the gas tracer.
The fundamental uncertainties in estimating the densities come from the limitation of the 3D extinction reconstruction, described in detail in section 6 of Edenhofer et al. (2024). We converted the extinction into nucleon density using the value from Bohlin et al. (1978), which recent observations of Hi emission and interstellar reddening indicate is too low by a factor of 1.5 Lenz et al. (2017). Variations from the diffuse environment at high Galactic latitude to the denser environment in the plane are also relevant, adding a factor of a few to the uncertainty. Recent observations also reveal spatial variability of $R_{V}$ characterized by a standard a standard deviation of around 0.25 (Zhang et al., 2023a).
The HOG analysis of synthetic line emission from a simulated MC presented in App. D indicates that the reconstructed Hi densities are roughly within a factor of two from the actual values. The spatial averaging in Eq. (3) results in smearing the density peaks across the area of a block, effectively limiting the maximum reconstructed density values. For CO, the comparison with the numerical model is limited by the extent of the synthetic diffuse CO emission, which is a complex chemical modeling problem beyond this work’s scope. However, the HOG method reproduces the CO density profile, although it over-estimates the CO density by a factor of a few. In either case, the limitations in the HOG reconstruction are below the variance across the studied region.
@@ 5.2 Kinetic energy
We estimated the average kinetic energy from streaming motions for each distance channel $q$ as
@@
where $X$ corresponds to the velocity tracer, either Hi or CO, $(v^{R19}_{\rm LOS})_{kq}$ is the expected LOS velocity from the Reid et al. (2019) Galactic rotation model for the Galactic plane position $k$ and the distance channel $q$, and $(v^{X}_{\rm LOS})_{kp^{*}q}$ is the LOS velocity from the line emission channel $p^{*}$ with the highest morphological correlation ($V_{\rm d}$) with the distance channel $q$. We calculated the effective volumetric density,
@@
using Eq. (3), the proton mass, $m_{p}$, and $\mu$ is the mean nucleon mass factor, which for solar metallicity is 1.402 (Asplund et al., 2009; Draine, 2010).
The $E_{\rm K}$ distribution is shown in the top-right panel of Fig. 14. We found that the average $E_{\rm K}$ is around 0.11 and 0.04 eV/cm3 for Hi and CO, respectively. These estimates are comparable to the 0.22 eV/cm3 obtained for the turbulent kinetic energy in the local ISM, obtained with a mean nucleon density $n$$=$ 30 cm-3 and velocity $v$$=$ 1 km s-1, or mean values $n_{\rm H}\right>$$=$ 1 cm-3 and $\sigma_{v}$$=$ 5.5 km s-1 (Draine, 2010). Given that we are only sampling one component of the velocity field, the actual $E_{\rm K}$ value can be up to a factor of three higher.
The $E_{\rm K}$ standard deviations are around 1.95 and 0.53 eV/cm3 for Hi and CO, respectively. These values, roughly a factor of four above the mean, suggest significant fluctuations of this quantity within the 2.5-kpc-diameter ISM parcel considered in this paper. The maximum $E_{\rm K}$ values are around 100 and 30 eV/cm3 for Hi In general, the highest $E_{\rm K}$ are associated with the highest $n^{\rm eff}$, suggesting that density is the dominant factor in our reconstruction of the kinetic energy distribution.
Figure 15 presents the $E_{\rm K}$ maps obtained from the Hi and CO emission. We found no evident similarity in the $E_{\rm K}$ distribution in the two gas tracers. However, for both tracers the highest $E_{\rm K}$ are found in the regions at $d$$<$ 300 pc.
The $E_{\rm K}$ radial profile, presented in the top panel of Fig. 16, shows values up to ten times over the mean for regions outside $d$$\approx$ 300 pc and up to 20 times for regions inside that range. The $E_{\rm K}$ peaks at $d$$<$ 300 pc do not seem to result from a concentration of high $V_{\rm d}$ due to a bias in the morphological correlation toward near distances, as discussed in App. C. However, the angular resolution in the 3D dust reconstruction may limit the mapping of $E_{\rm K}$ enhancements to further distances. One potential explanation for the $E_{\rm K}$ peak emerging at $d$$\approx$ 100 pc is the LSR location within the Local Bubble. We are observing the gas motions from a location near the center of that cavity so that the peak is at a similar radius in all directions. This is less likely to be the case for material that is further away from us.
@@
The $E_{\rm K}$ azimuthal profile, obtained by adding the energy along the LOS, is presented in the bottom panel of Fig. 16, It shows peaks toward $l$$\approx$ 40 to 90 for Hi and CO. This range covers the general directions along the axis of two large-scale structures identified in 3D dust reconstructions: the Split (Lallement et al., 2019) and the Radcliffe wave (Alves et al., 2020), shown in Fig. 3. Additional overdensities are found toward the fourth quadrant, roughly between $l$$\approx$ 300 to 340.
@@ 5.3 Momentum density and momentum
[…]
@@ 5.4 Mass flow rates
The $\dot{M}$ distribution, shown on the bottom-right panel of Fig. 14, is approximately centered around zero, implying that the LOS net mass flow for the studied region is close to zero. The $\dot{M}$ standard deviations are around 1.1 $\times$$10^{-3}$ and 0.4 $\times$$10^{-3}$ M/yr for Hi and CO, respectively. The azimuthal $\dot{M}$ profiles are presented in Fig 19. They correspond to the mean and the standard deviations of $\dot{M}$ along the LOS for Hi and CO. The mean mass flows, displayed in Fig 19, show net matter displacements of roughly couple $10^{-3}$ M/yr for lines of sight toward the first Galactic quadrant and 250 $<$$l$$<$ 360. This could be interpreted as residual velocity from the Galactic rotation if not because it has the opposite sign of the expected LOS motions indicated on the right-hand-side panel of Fig. 11. Further interpretation of these mass flows should consider that they may be a product of the LSR position and the boundaries of the 3D dust reconstruction.
@@ 6 Discussion
We found that roughly 67% and 15% of the 10$\times$ 10 distance channels in Fig. 5 show a significant morphological correlation, $\max(V_{\rm d})_{v}$$>$ 2.87, between the 3D dust and the Hi and the CO emission, respectively. Around 9% and 1% of the distance channels present $\max(V_{\rm d})_{v}$$<$ 0. Additional analysis of the $\max(V_{\rm d})_{v}$ distributions are presented in App. C.
@@
The lower correlation with CO can be explained by the lower volume filling factor of that tracer (Kwan & Sanders, 1986; Kalberla & Kerp, 2009). Clouds traced by CO are expected to be smaller and more concentrated than Hi structures. Hence, CO has fewer independent CO gradients for the HOG morphological comparison on the same sky area and at the common angular scale set by the derivative kernel size.
The lack of correlation in the rest of the channels can be attributed to several factors, which we separated into those related to the HOG method and those intrinsic to the dust and gas distribution in the local ISM. Among the former factors, we can distinguish two cases. First, there can be structures beyond 1.25 kpc with a morphological match with the emission within the input range. But even if they become available with a future 3D dust reconstruction, including them would not affect the results in Fig. 5 and only potentially fill a few gaps in the $lv$-diagrams presented in Fig. 6. Second, there can be line emission structures in channels beyond the $v_{\rm LOS}$ input range with a morphological match with the dust within the 3D reconstruction domain. We considered this scenario in App. A.3.3, where we reported the results for the input range $-120$$<$$v_{\rm LOS}$$<$ 120 km s-1. The expanded $v_{\rm LOS}$ range emphasizes higher streaming motion amplitudes at the inevitable price of an increased chance correlation. However, it does not patch the low $V_{\rm d}$ channels in Fig. 5.
@@
The low morphological correlation for certain portions of the studied volume can also be due to the intrinsic dust and gas distribution in the local ISM, which we consider in three main effects. First, there is low extinction in local ISM volumes, such as found in the regions around 600 $<$$d$$<$ 800  pc toward $-45$$<$$l$$<$ 30 and 1000 $<$$d$$<$ 1200  pc toward $250$$<$$l$$<$ 290, to mention a couple of examples. Second, line emission is absent across some LOS velocity ranges toward particular lines of sight. For example, toward the GSH 238+00+09 superbubble (Heiles, 1998), where there is very little CO, as illustrated in the right panel of Fig. 6. Identifying the exact physical conditions that produce the local ISM under-densities and a lack of emission toward particular regions is beyond the scope of this work. However, the maps in Fig. 5 and Fig. 6 provide a global context for identifying the conditions toward specific regions and their implications for the ISM dynamics at smaller scales.
@@
There is a third global factor that reduces the correlation between the 3D dust and the line emission. Observations of other ISM tracers, such as gamma-rays and the [Cii] and OH line emission, reveal that not all of the ISM material is traced by Hi and CO emission (see, for example, Grenier et al., 2005; Langer et al., 2014; Pineda et al., 2013; Busch et al., 2021). Including this material in the distance-velocity correlation is not straightforward. In the case of [Cii], there is no map with enough angular resolution and coverage to obtain significant spatial correlations using HOG. In the case of OH, the emission distribution involves the complex physical conditions that lead to the emission from that radical, which calls for a separate dedicated study (see, for example, Rugel et al., 2018; Dawson et al., 2022). We acknowledge that those and other line tracers should be accounted for in future reconstructions of the local ISM, adding richness to the results we present in this paper. Yet, our analysis has shown we have enough distance channels where the correlation between the dust distribution and the line emission is significant enough to identify prevalent LOS velocities for different portions of the local ISM, which we discuss in the following section.
@@ 6.2 Local ISM motions
We used the morphological correlation across distances and LOS velocities to assign a $v_{\rm LOS}$ to the distance channels in each of the 10$\times$ 10 regions in which we divided the studied volume. This assignment simplifies the velocity field by identifying a single value to ISM parcels with volumes between 102 to 105 pc3, corresponding to distance channels at the inner and outer edges of the 3D dust model. Thus, our velocity reconstruction represents the local ISM dynamics on spatial scales of tens of parsecs in the plane of the sky and around a parsec in the LOS direction.
@@
The maps of reconstructed LOS velocities shown in Fig. 11 indicate that the local LOS velocity field carries the imprint of the Galactic rotation pattern in both tracers, as inferred from the comparison with the expected $v_{\rm LOS}$ in the Reid et al. (2019) Galactic rotation model, shown in Fig. 10. The large-scale pattern of Hi and CO emission is deeply influenced by Galactic rotation, as shown in figure B.2 of HI4PI Collaboration et al. (2016) and figure 3 of Dame et al. (2001). However, the peculiar velocities of MCs within 3 kpc cannot be explained purely by Galactic rotation (see, for example, Stark & Brand, 1989).
@@
Figure 12 shows that the distribution of LOS streaming motions roughly centered around zero, within the uncertainties imposed by the velocity channel widths in the input data. The standard deviations of the rotation-subtracted LOS velocities are 10.8 and 6.6 km s-1 for Hi and CO, respectively. Previous studies have identified similar deviations from Galactic rotation. Clemens (1985) used the Massachusetts-Stony Brook Galactic plane CO survey to identify streaming motions with an amplitude of around 5 km s-1 on length scales larger than $0.22R_{\rm 0}$, where $R_{\rm 0}$ is the radius of the solar orbit around the Galactic center, so that $0.22R_{\rm 0}$ corresponds to a scale of between 1.87 and 2.2 kpc for the two values considered in that work. Brand & Blitz (1993) reported streaming motions with mean values of 12 km s-1 in their modeling of the spectro-photometric distances for a sample Hii regions and the LOS velocities for their associated MCs. The authors found a quadrupolar LOS velocity pattern, qualitatively comparable to those in Fig. 11, and report that the molecular gas is streaming past the LSR at about 3.8 km s-1, which are within a factor two we obtained for CO in the Solar vicinity. Our reported velocity dispersions in Hi are also comparable to those found in The Hi Nearby Galaxy Survey (THINGS; Walter et al., 2008; Schmidt et al., 2016). However, the physical conditions in their sample are not in general equivalent to those in the solar vicinity.